solve.
r^2-6r-16=0

Answer by ewatrrr(24785)


Hi,
solving
r^2-6r-16 = 0
factoring
(r-8)(r+2) = 0 Note:SUM of the inner product(-8r) and the outer product(2r) = -6r
(r-8)(r+2) = 0
(r-8)= 0
r = 8
(r+2) = 0
r = -2